A string of negligible mass going over clamped pulley of massm supports block of maas M the force on pulley by clamp

Grade:12

4 Answers

shubham khanna
8 years ago
F = ma

m = M
a = g (9.8)

F = 9.8M
Sneha Kharya
6 years ago
Force on pulley by the clamp = resultant of T = (M+m)gand mg acting along horizontal and vertical directions respectively F=√[{(M+m)g}^2+(mg)^2] =[√{(M+m)^2+m^2 }]g
Sohum
5 years ago
I am sorry to say that the answer given above is wrong. But there is only a small mistake that the horizontal force is notmg but it is Mg. And thus the answer chages to √[(M+m)^2+M^2]g. 
 
In future kindly read the question carefully before answering.
ankit singh
3 years ago
A string of negligible mass going over a clamped pulley of mass m supports a block of mass M as shown in the figure. The force on the pulley by the clamp is given by. =[√(M+m)2+m2]g .